OpenStudy (anonymous):

if r and s are the solutions of 2x^2+9x +3=0, what is the value of r^2 + s^2

OpenStudy (anonymous):

standard equation:\[ax ^{2} +bx +c =0\] and let r and s be two roots of this equation den r+s=(-b)/a r*s=c/a

OpenStudy (anonymous):

here r+s= -9/2 r*s=3/2 \[(r+s)^{2} =r ^{2} +s ^{2} +2timesr timess\]

OpenStudy (anonymous):

so 81/4 =r^2 +s^2 +2* (3/2) 81/4 = r^2 +s^2 +3 (81/4)-3 = r^2 +s^2 rest u can solve

OpenStudy (anonymous):

thx